OpenStudy (anonymous):

derivative of 1/(1+sec(x))^2

OpenStudy (anonymous):

\[\begin{align*} \frac{d}{dx}\left[\frac{1}{(1+\sec x)^2}\right]&=\frac{d}{dx}\left[(1+\sec x)^{-2}\right]\\\\ &=-2(1+\sec x)^{-3}\frac{d}{dx}\left[1+\sec x\right]\\\\ &=\cdots \end{align*}\]